if the roots of the equation x^2+3x+3=0 and x^2-x+lambda =0 are in the same ratio then the value of lambda is

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
first of all equation x^2+3x+3=0 has D =b2-4ac
A, B be the roots and the ratio be p so A=pB
then from equation one 
A+B=-b/2a
A.B= c/a 
and then putting A=pB  A and B were to be found and same thing was used for another equation. And lambda cound have been caculated
 
here (p+1)B=-3/2
p2B=3
calculte P here and b here and solve